P=α/β exp [-αz/Kbθ]
θ→temperature    P→pressure
Kb→Boltzmann constant      Z→distance
The dimensions of [α/β] is-

Plz explain what is exp and how α/β=P

 

Grade:11

1 Answers

Nicho priyatham
8 years ago
 
remember that exp(x) is called exponental function and
EXP( x) =e power x=ex =e^x
so the given equation is 
P= (α/β)e^[-αz/Kbθ] 
 
here u need to no 1 thing that is mathamatical functions like sinx, cosx, (all trignometric functions),logx, ex etc dont have dimensions  
 
 so in the equation P= (α/β)e^[-αz/Kbθ]   the term e^[-αz/Kbθ]  is                  diminsionless
 
so dimensions of P=dimensions of [α/β]
 PLZ APPROVE IF useful in case of any doubts drop it in ans box